# Star Formation Histories of Dwarf Lenticular Galaxies

### v1

@@ Abstract
We present the mean star formation histories (SFHs) of 148 dwarf lenticular galaxies (dS0s) derived from SDSS spectra. The SFHs of dS0s are characterized by multiple bursts of star formation, including an initial burst at a lookback time of Gyr for most galaxies. Stars formed during the first star-forming phase which ends at a lookback time of 6.3 Gyr primarily consist of old, metal-poor (Z=0.0004) stars , contributing to of the stellar mass and of the luminosity. The almost absence of extremely metal poor (Z=0.0001) stars seems to be due to pre-enrichment during the re-ionization era. Star formation gradually decreases during this initial period. In contrast, during the second period of star formation, there is an increase in star formation activity, peaking at a lookback time of 2.5 Gyr before declining again. Most moderately old stellar populations with intermediate metallicity were formed during this phase. We observe a strong dependence of SFHs on the mass and u-r color of dS0 galaxies but no significant dependence on morphological properties such as the presence or absence of outer spiral arms and nucleation. The star formation history of dS0 galaxies shares many similarities with that of dE galaxies, and many of them are believed to have originated from late-type galaxies.
galaxies.
@@
Dwarf lenticular galaxies (dS0s) are early-type dwarf galaxies whose morphology is similar to dwarf elliptical galaxies (dEs), but they possess a bulge/lens-like component in the inner regions. They were introduced by Sandage & Binggeli (1984) as a class of dwarf galaxies in the Virgo Cluster. Binggeli & Cameron (1991) conducted a detailed examination of the morphologies of dS0s, using 20 dS0s in the Virgo cluster. Their study revealed a variety of disc features present in the two-dimensional images of dS0s. The surface brightness profiles of dS0s resemble those of dEs, which are well fitted by exponential profiles or King profiles, displaying an excess of light in the central regions (Binggeli & Cameron, 1991, 1993). The similarity in surface brightness profiles of dEs and dS0s has raised questions regarding whether dS0s truly represent a distinct class of dwarf galaxies (e.g., Reyden et al., 1999), leading to some researchers treating dS0s as a subclass of dEs (e.g., Barazza et al., 2003).
@@
However, if we accept the definition of dE and dS0 galaxies as introduced by Sandage & Binggeli (1984), it appears reasonable to differentiate dS0s from dEs based on their surface brightness profiles, as dS0s exhibit multiple components—one for the inner regions and another for the outer regions. The surface brightness profile of the inner component is most accurately described by the Sérsic profile (Sérsic, 1968) while that of the outer component conforms to an exponential profile. The inner component is generally regarded as a lens (Buta, 2013), whereas the outer component is seen as a disc. Leveraging this characteristic of dS0s’ surface brightness, Aguerri et al. (2005) classified the early-type dwarfs in the Coma cluster into dEs and dS0s. They considered a galaxy to be a dE if a single component, the S’ersic profile, provided a good fit to the observed surface brightness profile. Conversely, they classified it as a dS0 if two components, Sérsic + exponential, were necessary to meet the constraint imposed on the residuals. Their method resulted in the identification of 29 dS0s and 55 dEs. A more comprehensive profile analysis was conducted by Janz et al. (2012, 2014) using GALFIT (Peng et al., 2010), a two-dimensional decomposition tool, on 121 early-type dwarfs in the Virgo cluster observed in the near-infrared -band. Only approximately one-third of the sample galaxies were found to possess a single component, while the majority of the 121 early-type dwarfs exhibited two or more components, including lenses and bars. Although they did not specify the morphological types of multi-component galaxies, a substantial proportion of multi-component dwarfs are likely to be dS0s.
@@
Despite the examples of dS0s provided by Binggeli & Cameron (1991) and the effectiveness of two-component models demonstrated by Aguerri et al. (2005), the ongoing debate revolves around whether dS0 should be considered a distinct, coequal category among early-type dwarfs. Many authors opt to refrain from explicitly designating dS0 as a separate class of early-type dwarf galaxies, instead employing dEs to encompass both dwarf elliptical galaxies and subsuming dSphs and dS0s within this category (e.g., Barazza et al., 2003). This tendency arises partly due to the limited number of dS0s in comparison to dEs, especially within the Virgo cluster, and the presence of disc features such as spiral arms, bars, and lenses observed in both dEs and dS0s (Jerjen et al., 2000, 2001; Barazza et al., 2002; Rijcke et al., 2003; Graham & Guzman, 2003; Graham et al., 2003; Ferrarese et al., 2006; Lisker et al., 2006; Lisker, 2009; Janz et al., 2012; Penny et al., 2014; Graham et al., 2017).
In light of these considerations, Lisker et al. (2006, 2007) made a distinction among early-type dwarf galaxies based on the presence or absence of embedded disc features, rather than isolating dS0 as a separate category. Galaxies devoid of disc features were categorized as dE, while those exhibiting such features were designated as dE(di).
The fraction of embedded disc features in dEs is no lower than that observed in dS0s (Seo & Ann, 2022). These embedded disc features are predominantly found in the brighter early-type dwarfs. The absence of disc features in the faint early-type dwarfs within the Virgo cluster aligns with the lack of such features in dSphs as well (Seo & Ann, 2022). While the dEs classified by Binggeli et al. (1985) encompass dSphs, which are generally faint early-type dwarfs, Ann et al. (2015) includes dSph as a class of early-type dwarfs in the morphology classification of galaxies in the local universe observed by the Sloan Digital Sky Survey (SDSS) (York et al., 2000).
It appears advisable to maintain the classification of dS0, even when employing dE(di) for other purposes, because the presence of embedded disc features is not unique to dS0s. Rather, the two-component nature of dS0s is an inherent characteristic of the dS0 type. Furthermore, the existence of spiral arms in the outer disc of certain dS0s represents a distinctive feature of this class. Ann et al. (2015) classified such cases as dS0, signifying peculiar dS0s.
@@
In the Virgo cluster, the count of dS0s is less than one-tenth of the count of dEs (Binggeli & Cameron, 1991; Kim et al., 2014). In contrast, in the local universe, the number of dS0s is approximately 60% of the number of dEs (Ann et al., 2015). One reason for the significant difference in the ratio of dEs to dS0s between the Virgo cluster and the local universe is that dSphs are classified as dEs in the VCC by Binggeli et al. (1985) and in the EVCC by Kim et al. (2014), whereas dSphs are categorized as a separate morphological type in the CVCG. Most of the faint dEs in the VCC and EVCC are classified as dSphs in the CVCG.
@@
The origin of dS0s is a subject of particular interest. Similar to dEs, there are two plausible scenarios for their formation: one related to their primordial origins and the other involving transformation processes. If dS0s are indeed transformed from late-type galaxies, the conversion of gas-rich disc galaxies into gas-poor early-type dwarfs likely involves two distinct phases. The initial phase entails gas removal, while the subsequent stage includes processes such as tidal heating and harassment. Given that a significant portion of dS0s are found in group and cluster environments (Makarov &Karachentsev, 2011; Ann, 2017), these transformation mechanisms may involve ram pressure stripping (Gunn & Gott, 1972) and galaxy harassment (Moore et al., 1996, 1998).
@@
In contrast, Boselli et al. (2008) and Steyrleithner et al. (2020) have reported that the transformation of late-type galaxies into early-type dwarfs can be achieved solely through ram pressure stripping, while Chilingarian et al. (2009) suggested the possibility of major/minor mergers playing a role. The morphological transformation from late-type galaxies to early-type dwarfs is feasible because ram pressure not only removes cold gas but also compresses it, leading to star formation. The grand-design spiral arms observed in unsharp masked images of some early-type galaxies (e.g., Lisker et al., 2006) are believed to form in the gas compressed by ram pressure. The disc features, including arms, bars, lenses, and clumps, seen in the residual images of early-type dwarfs, especially dS0s, could result from star formation driven by ram pressure. Ram pressure has the effect of heating a thin stellar disc, turning it into a thicker one by removing the gas potential (Smith et al., 2012). Additionally, gas clouds pushed out of the disc by ram pressure but still gravitationally bound may fall back in and contribute to star formation in the thick disc (Steyrleithner et al., 2020; Boselli et al., 2022).
@@
Conversely, the hypothesis of a primordial origin for dS0s remains a viable one, supported by the fact that approximately 5% of them exist as isolated galaxies, unaffected by environmental factors that could potentially alter their morphology. However, recent analyses of the star formation histories (SFHs) of dEs and dSphs (Seo & Ann, 2023), derived from SDSS spectra, have revealed that a significant fraction of dEs cannot be considered primordial objects, as they exhibit active periods of star formation around a lookback time of 2.5 Gyr, whereas the majority of dSphs are believed to be primordial objects. Since dS0s share many similarities with dEs, it is reasonable to expect that the SFHs of dS0s are similar to those of dEs.
@@
There have been relatively few studies on the SFHs of dS0s, primarily due to the scarcity of dS0s within the Local Group (LG), where SFHs of dwarf galaxies can be derived using the Color-Magnitude Diagram (CMD) method (Tolstoy et al., 2009). The primary objective of this paper is to investigate the origin of dS0 galaxies by analyzing how their SFHs. It is well-established that stellar mass plays a pivotal role in driving star formation in galaxies, resulting in the earlier formation of stars in more massive galaxies compared to less massive ones (Cowie et al., 1996; Gavazzi et al., 2006; De Lucia et al., 2007; Sánchcez-BLázquez, 2009). The relationship between star formation rates (SFRs) and stellar mass is also well-documented (Brinchmann et al., 2004; Daddi et al., 2007; Noeske et al., 1997; Salim et al., 2007; Rodighiero et al., 2011; Sobral et al., 2014; Speagle et al., 2014).
@@
The influence of the environment on the SFHs of galaxies has been studied extensively for bright galaxies (Spitzer & Baade, 1951; Oemler, 1974; Dressler, 1980; Kauffmann et al., 2004; Wang, 2022; Pérez-Millán et al., 2023), as well as for early-type dwarfs, specifically dSphs and dEs (Seo & Ann, 2023). In this research, we focus on the SFHs of dS0s, paying special attention to the dependence of the cumulative star formation histories (cSFHs) on the physical and morphological properties of dS0s, as well as their environemnt. To achieve this, we derive the stellar populations by applying the population synthesis code STARLIGHT (Cid Fernandes et al., 2005) to the spectra of these galaxies observed by the SDSS.
@@
The structure of this paper is organized as follows. In Section 2, we outline the process of selecting sample galaxies and provide a brief overview of the method for analyzing spectra using STARLIGHT. Section 3 presents the SFHs of dS0s, focusing on luminosity and mass fractions. In Section 4, we present the cSFHs obtained from our analysis. Finally, in Section 5, we conclude the study by discussing our findings
@@ 2.1 Data
We utilized the SDSS spectra of 148 dS0 galaxies, all of which are listed in the CVCG. While the CVCG initially includes 154 dS0 galaxies, we were able to obtain spectra for 148 dS0s from the SDSS Data Release 7 (DR7). It’s worth noting that the CVCG encompasses a total of 5,638 galaxies situated in the local universe. Furthermore, the CVCG is nearly complete for galaxies brighter than in the regions covered by the SDSS. The detailed classification of dS0 galaxies was provided by Ann et al. (2015), who distinguished dS0s with spiral arm features as peculiar dS0s, and also differentiated between nucleated and non-nucleated dS0s. Among the 148 dS0 galaxies, 30 exhibit spiral arms in the outer regions, constituting roughly 20% of the dS0 population. Concerning nucleation, the number of nucleated dS0 galaxies is comparable to that of non-nucleated dS0s.
Figure 1 presents color images of four sample galaxies, with two belonging to dS0s without spiral arm features (upper row) and the other two to dS0s with spiral arm features in the outer regions (lower row). In the right panel of each pair, we show nucleated dS0s, while the left panel displays non-nucleated dS0s. A notable observation from these images is the presence of a lens-like component in the inner regions of the dS0s, which is a distinctive feature.
As illustrated in Figure 2, dS0s with spiral arm features appear slightly fainter compared to those without such features. It’s worth mentioning that nearly all dS0s with spiral arm features are fainter than . Furthermore, the majority of dS0s are situated along the red sequence (Strateva et al., 2001), although there are some notable deviations. The significant dispersion in the colors among dS0s is likely attributed to the relatively large photometric errors in -band magnitudes, especially for faint galaxies.
The SDSS spectra were captured using fibers with a diameter of 3 arcsec, positioned at the focal plane of the 2.5-meter telescope located at the Apache Point Observatory. The SDSS spectrograph incorporates 320 fibers, and the exposures lasted for 45 minutes or more, ensuring a fiducial signal-to-noise ratio. The spectra encompass a wavelength range spanning from 3800 to 9200 Å, with an average spectral resolution of . The wavelength and flux values have been calibrated using the pipeline developed by the SDSS team. We have also incorporated observational data from Ann et al. (2015), which includes information such as distance, luminosity (M), color (), morphological type, as well as coordinates and redshift.
@@ 2.2 Method
STARLIGHT fits the observed spectrum by generating a model spectrum derived from population synthesis models based on simple stellar populations (SSPs) from Bruzual & Charlot (2003). These models incorporate six metallicities (Z = 0.0001, 0.0004, 0.004, 0.008, 0.02, and 0.05) and ages ranging from to years to represent the spectral evolution of SSPs. The stellar populations encompass stars with lower and upper mass cut-offs of and . Bruzual & Charlot (2003) utilized basic stellar evolutionary tracks from the Padova groups, complemented by the Geneva groups’ data, along with the STELIB library (Le Borgne et al., 2003). In the STARLIGHT framework, the model spectrum is represented as
where is the th SSP spectrum normalized at , , is the synthetic flux at the normalization wavelength , and is the fractional contribution of the SSP for th population that has age and metallicity Z.The stellar motion projected on the line-of-sight is modeled by a Gaussian distribution (G) centered on the galaxy radial velocity with velocity dispersion . Extinction due to foreground dust is taken into account using the V-band extinction A. The best fitting model is determined by selecting a model that minimizes the where is observed spectrum and is the inverse of error applied (see Cid Fernandes et al. (2004) for a detailed description).
@@
Figure 3 shows an example of population systhesis carried out using STARLIGHT. The model spectrum well represents the details of the observed spectrum of IC 3540 which is a dwarf lenticular galaxy in the Virgo cluster. The robustness of the stellar populations derived by STARLIGHT has been examined in previous works (Magris et al., 2015; Cid Fernandes, 2018; Seo & Ann, 2023). To assess its reliability, Seo & Ann (2023) conducted mock tests using STARLIGHT, where they generated mock spectra from model fluxes perturbed by Gaussian noise levels proportional to . They observed that ages and metallicities of mock galaxies were well reproduced, except for spectra with = 5. Here we present a comparision of the mean ages and metallicities derived from STARLIGHT and pPXF (Cappellari & Emsellem, 2004). The mean ages and metallicities are calculated using luminosity-weighted values through the following equations,
where is the fractional luminosity contribution of -th population and and are its stellar age and metallicity, respectively. As shown in Figure 4, the two data sets show reasonable agreement which enforces the robustness of the methods applied if we treat them statistically.

### v2

@@ Abstract
Star formation histories (SFHs) are essential for understanding galaxy formation and evolution. We present the mean SFHs of 148 dwarf lenticular galaxies (dS0s) derived from the Sloan Digital Sky Survey (SDSS) spectra. These SFHs reveal two distinct periods of active star formation. The first period, concluding approximately 6 Gyr ago, witnessed the formation of predominantly old, metal-poor (Z=0.0004) stars, accounting for approximately 60% of the stellar mass and 30% of the luminosity. The scarcity of extremely metal-poor (Z=0.0001) stars suggests pre-enrichment during the re-ionization era. Star formation gradually waned during this initial phase. In contrast, the second period, ending around 1 Gyr ago, exhibited a peak in the middle of the period, contributing to the formation of moderately old stellar populations with intermediate metallicity. The SFHs of dS0 galaxies show a clear dependence on stellar mass, with more massive dS0s forming stars earlier. But, we find no significant correlation with morphological properties such as outer spiral arms and nucleation. The SFHs of dS0 galaxies share many similarities with those of dE galaxies, suggesting a common origin, mostly not primordial.
@@
Dwarf lenticular galaxies (dS0s) are early-type dwarf galaxies whose morphology is similar to dwarf elliptical galaxies (dEs), but they possess a bulge/lens-like component in the inner regions. They were introduced by Sandage & Binggeli (1984) as a class of dwarf galaxies in the Virgo Cluster. Binggeli & Cameron (1991) conducted a detailed examination of the morphologies of dS0s, using 20 dS0s in the Virgo cluster. Their study revealed a variety of disc features present in the two-dimensional images of dS0s. The surface brightness profiles of dS0s resemble those of dEs, which are well fitted by exponential profiles or King profiles, with an excess of light in the central regions (Binggeli & Cameron, 1991, 1993). The similarity in surface brightness profiles of dEs and dS0s has raised questions regarding whether dS0s truly represent a distinct class of dwarf galaxies (e.g., Reyden et al., 1999), leading to some researchers treating dS0s as a subclass of dEs (e.g., Barazza et al., 2003).
@@
However, if we accept the definition of dE and dS0 galaxies as introduced by Sandage & Binggeli (1984), it seems reasonable to differentiate dS0s from dEs based on their surface brightness profiles, as dS0s exhibit multiple components with one for the inner regions and another for the outer regions. The surface brightness profile of the inner component is most accurately described by the Sérsic profile (Sérsic, 1968), while that of the outer component conforms to an exponential profile. The inner component is generally regarded as a lens (Buta, 2013), whereas the outer component is seen as a disc. Leveraging this characteristic of dS0s’ surface brightness, Aguerri et al. (2005) classified the early-type dwarfs in the Coma cluster into dEs and dS0s while Janz et al. (2012, 2014) classified the early-type dwarfs in the Virgo cluster using GALFIT (Peng et al., 2010), a two-dimensional decomposition tool. Their classification revealed significantly smaller dS0/dE ratios in the Coma cluster compared to the Virgo cluster. The reason for the higher fraction of dS0s in the Virgo cluster seems to be the better resolution owing to the shorter distance to the Virgo cluster. Additionally, the use of two-dimensional images rather than one-dimensional surface brightness profiles employed by Aguerri et al. (2005) may have contributed to this classification difference.
@@
Despite the examples of dS0s provided by Binggeli & Cameron (1991) and the effectiveness of two-component models demonstrated by Aguerri et al. (2005), the ongoing debate revolves around whether dS0 should be considered a distinct, coequal category among early-type dwarfs. Many researchers prefer to classify all dwarf elliptical-like galaxies, including dwarf spheroidals (dSphs) and dwarf lenticulars (dS0s), under the broader category of dE galaxies (e.g., Barazza et al., 2003). This tendency arises partly due to the limited number of dS0s in comparison to dEs, especially within the Virgo cluster, and the presence of disc features such as spiral arms, bars, and lenses observed in both dEs and dS0s (Jerjen et al., 2000, 2001; Barazza et al., 2002; Rijcke et al., 2003; Graham & Guzman, 2003; Graham et al., 2003; Ferrarese et al., 2006; Lisker et al., 2006; Lisker, 2009; Janz et al., 2012; Penny et al., 2014; Graham et al., 2017).
@@
The origin of dS0s is a subject of particular interest. Similar to dEs, there are two plausible scenarios for their formation: one related to their primordial origins and the other involving transformation processes. If dS0s are indeed transformed from late-type galaxies, the conversion of gas-rich disc galaxies into gas-poor early-type dwarfs likely involves two distinct phases. The initial phase entails gas removal, while the subsequent stage includes processes such as tidal heating and harassment. Given that a significant portion of dS0s are found in group and cluster environments (Makarov &Karachentsev, 2011), these transformation mechanisms may involve ram pressure stripping (Gunn & Gott, 1972) and galaxy harassment (Moore et al., 1996, 1998).
@@
In contrast, Boselli et al. (2008) and Steyrleithner et al. (2020) have reported that the transformation of late-type galaxies into early-type dwarfs can be achieved solely through ram pressure stripping, while Chilingarian et al. (2009) suggested the possibility of major/minor mergers playing a role. The morphological transformation from late-type galaxies to early-type dwarfs is feasible because ram pressure not only removes cold gas but also compresses it, leading to star formation. The grand-design spiral arms observed in unsharp masked images of some early-type galaxies (e.g., Lisker et al., 2006) are believed to form in the gas compressed by ram pressure. The disc features, including arms, bars, lenses, and clumps, seen in the residual images of early-type dwarfs, especially dS0s, could result from star formation driven by ram pressure. Ram pressure has the effect of heating a thin stellar disc, turning it into a thicker one by removing the gas potential (Smith et al., 2012). Additionally, gas clouds pushed out of the disc by ram pressure can remain gravitationally bound, eventually falling back and fueling star formation in the thick disc (Steyrleithner et al., 2020; Boselli et al., 2022).
@@
Conversely, the hypothesis of a primordial origin for dS0s remains a viable one, supported by the fact that approximately 7% of them exist as isolated galaxies (Ann & Seo, 2024), unaffected by environmental factors that could potentially alter their morphology. However, recent analyses of the star formation histories (SFHs) of dEs and dSphs (Seo & Ann, 2023), derived from the SDSS spectra, have revealed that a significant fraction of dEs cannot be considered primordial objects, as they exhibit active periods of star formation around 2.5 Gyr ago, whereas the majority of dSphs are believed to be primordial objects. Since dS0s share many similarities with dEs, it is reasonable to expect that the SFHs of dS0s are similar to those of dEs.
@@
There have been relatively few studies on the SFHs of dS0s, primarily due to the scarcity of dS0s within the Local Group (LG), where SFHs of dwarf galaxies can be derived using the Color-Magnitude Diagram (CMD) method (Tolstoy et al., 2009). The primary objective of this paper is to investigate the origin of dS0 galaxies by analyzing their SFHs. It is well-established that stellar mass plays a pivotal role in driving star formation in galaxies, resulting in the earlier formation of stars in more massive galaxies compared to less massive ones (Cowie et al., 1996; Gavazzi et al., 2006; De Lucia et al., 2007; Sánchcez-BLázquez, 2009). The relationship between star formation rates (SFRs) and stellar mass is also well-documented (Brinchmann et al., 2004; Daddi et al., 2007; Noeske et al., 2007; Salim et al., 2007; Rodighiero et al., 2011; Sobral et al., 2014; Speagle et al., 2014).
@@
The influence of the environment on the SFHs of galaxies has been studied extensively for bright galaxies (Spitzer & Baade, 1951; Oemler, 1974; Dressler, 1980; Kauffmann et al., 2004; Wang, 2022; Pérez-Millán et al., 2023), as well as for early-type dwarfs, specifically dSphs and dEs (Seo & Ann, 2023). In this research, we focus on the SFHs of dS0s, paying special attention to the dependence of the cumulative star formation histories (cSFHs) on the physical and morphological properties of dS0s, as well as their environment. To achieve this, we derive the stellar populations by applying the population synthesis code STARLIGHT (Cid Fernandes et al., 2005) to the spectra of these galaxies observed by the SDSS.
@@
The structure is organized as follows. In Section 2, we outline the process of selecting sample galaxies and provide a brief overview of the method for analyzing spectra using STARLIGHT. Section 3 presents the SFHs of dS0s, focusing on luminosity and mass fractions. In Section 4, we present the cSFHs obtained from our analysis. Finally, in Section 5, we conclude the study by discussing our findings.
@@ 2.1 Data
We utilized the SDSS spectra of 148 dS0 galaxies, all of which are listed in the CVCG. While the CVCG initially includes 154 dS0 galaxies, we were able to obtain spectra for 148 dS0s from the SDSS Data Release 7 (DR7). It’s worth noting that the CVCG encompasses a total of 5,638 galaxies situated in the local universe. Furthermore, the CVCG is nearly complete for galaxies brighter than in the regions covered by the SDSS. The signal-to-noise ratio (S/N) of their spectra mostly falls within the range of 10 to 30. The detailed classification of dS0 galaxies was provided by Ann et al. (2015), who distinguished dS0s with spiral arm features as peculiar dS0s, and also differentiated between nucleated and non-nucleated dS0s. Among the 148 dS0 galaxies, 30 exhibit spiral arms in the outer regions, constituting roughly 20% of the dS0 population. Concerning nucleation, the number of nucleated dS0 galaxies is comparable to that of non-nucleated dS0s.
Figure 1 presents color images of four dS0 galaxies: two without spiral arms (upper row) and two with spiral arms (lower row). The right panels show nucleated dS0s, while the left panels display non-nucleated dS0s. A notable observation from these images is the presence of a lens-like component in the inner regions of the dS0s, a distinctive feature.
As illustrated in Figure 2 where we can see the distribution of all CVCG galaxies in the versus color-magnitude diagram, dS0s with spiral arm features appear slightly fainter compared to those without such features. It’s worth mentioning that nearly all dS0s with spiral arm features are fainter than . Furthermore, the majority of dS0s are situated along the red sequence (Strateva et al., 2001), if we consider the photometric error of color as . The error of is mostly due to distance error () and we assumed to derive the error of in Figure 2.
The SDSS spectra were captured using fibers with a diameter of 3 arcsec, positioned at the focal plane of the 2.5-meter telescope located at the Apache Point Observatory. The SDSS spectrograph incorporates 320 fibers, and the exposures lasted for 45 minutes or more, ensuring a fiducial signal-to-noise ratio. The spectra encompass a wavelength range spanning from 3800 to 9200 Å, with an average spectral resolution of . The wavelength and flux values have been calibrated using the pipeline developed by the SDSS team. We have also incorporated observational data from Ann et al. (2015), which includes information such as distance, luminosity (Mr), color (), morphological type, as well as coordinates and redshift.
@@ 2.2 Method
STARLIGHT fits the observed spectrum by generating a model spectrum derived from population synthesis models based on simple stellar populations (SSPs) from Bruzual & Charlot (2003). These models incorporate six metallicities (Z = 0.0001, 0.0004, 0.004, 0.008, 0.02, and 0.05) and 25 age bins ranging from log (age)= 6 to 10.215 to represent the spectral evolution of SSPs. The stellar populations encompass stars with lower and upper mass cut-offs of and . Bruzual & Charlot (2003) utilized basic stellar evolutionary tracks from the Padova groups, complemented by the Geneva groups’ data, along with the STELIB library (Le Borgne et al., 2003).
In the STARLIGHT framework, the model spectrum is represented as
where is the th SSP spectrum normalized at , , is the synthetic flux at the normalization wavelength , and is the fractional contribution of the SSP for th population that has age and metallicity Zj.The stellar motion projected on the line-of-sight is modeled by a Gaussian distribution (G) centered on the galaxy radial velocity with velocity dispersion . Extinction due to foreground dust is taken into account using the V-band extinction AV and the reddening law of Cardelli et al. (1980). We calculated AV using E(B-V) obtained from the dustmaps (Schlegel et al., 1998). The best fitting model is determined by selecting a model that minimizes the where is observed spectrum and is the inverse of error applied (see Cid Fernandes et al. (2004) for a detailed description).
@@
2.3 Uncertainties
Figure 3 shows an example of population synthesis carried out using STARLIGHT. The model spectrum closely matches the details of the observed spectrum of IC 3540, a dwarf lenticular galaxy in the Virgo cluster. However, it is not a mathematical solution but a statistical solution, implying uncertainties in the derived physical parameters, such as stellar mass, mean age, and mean metallicity, are inevitable. Several factors introduce errors in the products of population synthesis models.
Noise in the observed spectrum is the primary source of errors. Seo & Ann (2023) performed a mock test to examine the effect of random noise on the mean age and metallicity of 434 early-type dwarfs. They showed correlations between the mean metallicities and ages of stellar populations derived from the mock spectra and those of models, alongside the root mean square (rms) error of the mean metallicities. We used this data (their Figure 4) to calculate the rms errors of the mean ages derived from STARLIGHT. They range from 0.1 dex for spectra with signal-to-noise ratio () of 30 to 0.2 dex for spectra with . Our estimate of uncertainty in the mean age agrees with that of Cid Fernandes et al. (2005). Similar uncertainty was also reported by Magris et al. (2015), who noted that the uncertainty in log(age) for STARLIGHT is roughly the same as that of other nonparametric population synthesis methods. The lower limit of uncertainty in log (age) is set by the grid of isochrones in the base models. However, uncertainty of dex in log (age) is internal. There are biases resulting from the population synthesis models, which use different sets of SSPs as well as different numerical methods. The model dependence of the derived physical parameters was discussed by Cid Fernandes et al. (2014), who showed 0.2 dex difference in mass, mean age, and metallicity among different base models.
Here we present a comparision of the mean ages and metallicities derived from STARLIGHT and pPXF (Cappellari & Emsellem, 2004). The mean ages and metallicities are calculated using luminosity-weighted values through the following equations,
where is the fractional luminosity contribution of -th population and and are its stellar age and metallicity, respectively. As shown in Figure 4, the two datasets exhibit a correlation with significant scatter between them. The rms errors for both parameters are approximately dex. Moreover, the mean metallicities (upper panel) derived from pPXF deviate systematically from the one-to-one relation by -0.17 dex. However, the large rms errors are expected because, as described above, the rms error due to noise in the observed spectra amounts to approximately dex, and the rms error originating from the differences in the base models and calculation methods is about dex (Cid Fernandes et al., 2014; Magris et al., 2015). Thus, it appears better to treat higher-order physical parameter such as SFHs statistically, although physical parameters of individual galaxies still provide important information about their SFHs.